OpenStudy (anonymous):

Find the zeros of the function f(x)=2x^2-16x+32 by factoring.

hero (hero):

Set f(x) = 0: 2x^2-16x+32 = 0 2(x^2-8x+16) = 0 x^2-8x + 16 = 0 x^2-4x-4x+16 = 0 x(x-4)-4(x-4) = 0 (x-4)(x-4) = 0 (x-4)^2 = 0 x - 4 =0 x = 4

OpenStudy (anonymous):

Thanks for writing out how you did it... it really helps

hero (hero):

You're most welcome

hero (hero):

Appreciate you Sheg